if force F and velocity v is related to time t & mass m, acceleration a according to the equation


F/a=tx my 


then the value of x=


1
Expert's answer
2021-08-29T16:56:19-0400

The units of accelration are:


"[a] = m\/s^2"

of force:


"[F] = kg\cdot m/s^2"

of time "[t] = s" and mass "[m] = kg". Thus, obtain:


"\dfrac{[F]}{[a]} = [t]^x[m]^y\\
\dfrac{kg\cdot m\cdot s^2}{s^2\cdot m} = kg^x\cdot s^y\\
kg = kg^x\cdot s^y"

Comparing the sides of the las equation, obtain:


"x = 1, y = 0"

Answer. "x = 1, y = 0."